# Radial evolution of Alfvén wave Parametric Decay Instability in the near-Sun solar wind: Effects of Temperature Anisotropy

### 24 Apr 2026 v1

@@ Abstract
Parametric decay instability (PDI) of Alfvén wave is thought to play an important role in the dissipation of the large-amplitude Alfvén waves and in the heating of magnetized plasmas. Temperature anisotropy is frequently observed by spacecrafts, including Parker Solar Probe (PSP), in the near-Sun solar wind, yet its impact on PDI in the near-Sun solar wind has been understudied. We calculate the maximum growth rates of PDI, $\gamma_{\max}/\omega_{0}$, where $\omega_{0}$ is the frequency of the parent wave, by solving the linear dispersion relation of Chew-Goldberger-Low (CGL) equations under several expanding background models. To assess the effect of temperature anisotropy, the growth rate is compared with that derived from ideal magnetohydrodynamics (MHD). From $R_{0}$ ($\sim 1.02R_{\odot}$) to $30R_{0}$, we consider three expansion cases: (i) spherically symmetric adiabatic expansion with constant wind speed, (ii) Multi-source observation- and model-constrained expansion, and (iii) a PSP-constrained profile of $(\beta_{\parallel},\xi)$, where $\beta_{\parallel}=8\pi p_{\parallel 0}/B_{0}^{2}$ is the parallel plasma beta and $\xi=T_{\perp 0}/T_{\parallel 0}$ is the temperature anisotropy, that includes Parker-spiral effects. We find that temperature anisotropy increases $\gamma_{\max}/\omega_{0}$ for $\beta\lesssim 0.1$ in the near-Sun solar wind: in the case of (iii), temperature anisotropy with $T_{\perp 0}>T_{\parallel 0}$ increases $\gamma_{\max}/\omega_{0}$ by factors of $\sim 1.5$ over $R\simeq 1$$15\,R_{0}$, whereas temperature anisotropy with $T_{\parallel 0}>T_{\perp 0}$ decreases $\gamma_{\max}/\omega_{0}$ at larger $R$. Our results suggest that the temperature anisotropy plays an important role in the onset of PDI even in low-$\beta$ regimes, such as the near-Sun solar wind.
@@ II.2.1 Case 1: Spherically symmetric adiabatic expansion with constant wind speed
In terms of the temperature anisotropy, we apply the double-adiabatic closures such as equations of (2.11) and (2.12) to the background. Then we get
@@ IV.2 The effect of the perturbed temperature anisotropy
The differences between the maximum growth rates obtained from the isotropic and anisotropic dispersion relations arise from both the background temperature anisotropy and the anisotropy of the perturbations ($\delta T_{\perp}\neq\delta T_{\parallel}$). Even if we impose a background temperature anisotropy of unity in the CGL-based dispersion relation, (2.13), the resulting dispersion relation remains completely different from the isotropic (ideal-MHD) dispersion relation, (2.6). The reason why the dispersion relations differ even under identical background conditions is that the pressure perturbations—i.e., the linearized pressure tensor $\delta{P}$—are not actually the same in the ideal-MHD and CGL models. In other words, because the two linearized sets of governing equations are not equivalent, the dispersion relations derived from them are also different. However, it can be shown that the two linearized systems become identical if, in addition to setting $\xi=1$, we further impose $\delta T_{\parallel}=\delta T_{\perp}$, provided that the polytropic index on the MHD side is assumed to be $5/3$ (Appendix B).
@@ V Conclusion and Summary
    Background temperature anisotropy can significantly enhance PDI even in low-$\beta$ near-Sun conditions. In case 3, where $\beta_{\parallel}\lesssim 0.1$ in the inner heliosphere, the CGL model yields a maximum growth rate larger than the ideal-MHD value by a factor of $\sim 1.5$ over $R/R_{0}\sim 2$$12$ when the background anisotropy is perpendicular-temperature dominated. Moreover, in case 3, temperature anisotropy with $T_{\parallel}>T_{\perp}$ reduces the maximum growth rate, thereby limiting the radial extent of the PDI-unstable region to $R\lesssim 30R_{0}$. Other effects—such as solar-wind expansion, background turbulence, and Landau damping—are not considered in this study, so our estimated radial range for PDI occurrence could be confined to even smaller heliocentric distances. Our result demonstrates that anisotropy-driven enhancement remains effective in the parameter regime most relevant to PSP observations, and therefore temperature anisotropy should be treated as a non-negligible control parameter when assessing where PDI can operate in the near-Sun wind.
@@
    In the adiabatic case, the radial evolution of $\gamma_{\max}/\omega_{0}$ differs qualitatively between the isotropic and anisotropic closures. Under the adiabatic assumption, ideal MHD predicts that $\gamma_{\max}/\omega_{0}$ increases with heliocentric distance in the low-$\beta$ regime, with an approximate power-law scaling that approaches the classical ${\gamma}_{\max}/\omega_{0}\propto R^{1/3}$ behavior in the smaller-amplitude limit. In contrast, the CGL model does not exhibit a monotonic increase with $R$. Instead, ${\gamma}_{\max}/\omega_{0}$ can turn over and decrease beyond a certain distance. The parameter scans in Section IV.1 suggest that this difference is primarily related to the different $\beta_{\parallel}$-dependence implied by the two closures: in the low-$\beta$ regime, the PDI growth rates in both models are nearly identical even when $\xi=10$ is imposed, whereas the direct $\xi$-dependence itself remains weak over a broad range. Because the adiabatic closures imply different radial scalings, $\beta\propto R^{2/3}$ in MHD and $\beta_{\parallel}\propto R^{2}$ in CGL, the more rapid increase of $\beta_{\parallel}$ in the CGL closure provides a plausible explanation for the different radial behaviors found in case 1.
@@
    Perturbed temperature anisotropy reduces ${\gamma}_{\max}/\omega_{0}$. Even when the background anisotropy is set to unity ($\xi=1$), the CGL dispersion relation does not reduce to the ideal-MHD form because the linearized pressure responses differ unless one further imposes $\delta T_{\perp}=\delta T_{\parallel}$ (for $n=5/3$). By comparing ${\gamma}_{\max}/\omega_{0}$ from CGL at $\xi=1$ with the ideal-MHD result, we quantified the impact of the perturbed temperature anisotropy. The resulting maps show that ${\gamma}_{\max}/\omega_{0}$ is generally smaller in CGL than in MHD, and the difference grows with increasing $\hat{B}_{\perp}$ and $\beta$. This implies that allowing $\delta T_{\perp}\neq\delta T_{\parallel}$ tends to suppress PDI.
These results have direct implications for PDI-based interpretations and models of the near-Sun solar wind. Because PDI produces compressive fluctuations and counter-propagating Alfvénic components, any temperature-anisotropy-driven modulation of ${\gamma}_{\max}/\omega_{0}$ can influence the generation of density fluctuations, the effective reflection rate, and the subsequent development of Alfvén-wave turbulence. Our findings therefore motivate incorporating anisotropic thermodynamics when estimating PDI activity in inner-heliosphere applications, including reduced models and global simulations that use PDI as an effective mechanism for controlling turbulence.

### 27 Apr 2026 v2

@@ Abstract
Parametric decay instability (PDI) of Alfvén wave is thought to play an important role in the dissipation of the large-amplitude Alfvén waves and in the heating of magnetized plasmas. Temperature anisotropy is frequently observed by spacecrafts, including Parker Solar Probe (PSP), in the near-Sun solar wind, yet its impact on PDI in the near-Sun solar wind has been understudied. We calculate the maximum growth rates of PDI, $\gamma_{\max}/\omega_{0}$, where $\omega_{0}$ is the frequency of the parent wave, by solving the linear dispersion relation of Chew-Goldberger-Low (CGL) equations under several expanding background models. To assess the effect of temperature anisotropy, the growth rate is compared with that derived from ideal magnetohydrodynamics (MHD). From $R_{0}$ ($1.02R_{\odot}$) to $30R_{0}$, we consider three expansion cases: (i) spherically symmetric adiabatic expansion with constant wind speed, (ii) Multi-source observation- and model-constrained expansion, and (iii) a PSP-constrained profile of $(\beta_{\parallel},\xi)$, where $\beta_{\parallel}=8\pi p_{\parallel 0}/B_{0}^{2}$ is the parallel plasma beta and $\xi=T_{\perp 0}/T_{\parallel 0}$ is the temperature anisotropy, that includes Parker-spiral effects. We find that temperature anisotropy increases $\gamma_{\max}/\omega_{0}$ for $\beta\lesssim 0.1$ in the near-Sun solar wind: in the case of (iii), temperature anisotropy with $T_{\perp 0}>T_{\parallel 0}$ increases $\gamma_{\max}/\omega_{0}$ by factors of $\sim 1.5$ over $R\simeq 1$$10\,R_{0}$, whereas temperature anisotropy with $T_{\parallel 0}>T_{\perp 0}$ decreases $\gamma_{\max}/\omega_{0}$ at larger $R$. Our results suggest that the temperature anisotropy plays an important role in the onset of PDI even in low-$\beta$ regimes, such as the near-Sun solar wind.
@@ II.2.1 Case 1: Spherically symmetric adiabatic expansion with constant wind speed
In terms of the temperature anisotropy, we apply the double-adiabatic closures such as equations of (II.10) and (II.11) to the background. Then we get
@@ IV.2 The effect of the perturbed temperature anisotropy
The differences between the maximum growth rates obtained from the isotropic and anisotropic dispersion relations arise from both the background temperature anisotropy and the anisotropy of the perturbations ($\delta T_{\perp}\neq\delta T_{\parallel}$). Even if we impose a background temperature anisotropy of unity in the CGL-based dispersion relation, (II.12), the resulting dispersion relation remains completely different from the isotropic (ideal-MHD) dispersion relation, (II.6). The reason why the dispersion relations differ even under identical background conditions is that the pressure perturbations—i.e., the linearized pressure tensor $\delta{P}$—are not actually the same in the ideal-MHD and CGL models. In other words, because the two linearized sets of governing equations are not equivalent, the dispersion relations derived from them are also different. However, it can be shown that the two linearized systems become identical if, in addition to setting $\xi=1$, we further impose $\delta T_{\parallel}=\delta T_{\perp}$, provided that the polytropic index on the MHD side is assumed to be $5/3$ (Appendix B).
@@ V Conclusion and Summary
    Background temperature anisotropy can significantly enhance PDI even in low-$\beta$ near-Sun conditions. In case 3, where $\beta_{\parallel}\lesssim 0.1$ in the inner heliosphere, the CGL model yields a maximum growth rate larger than the ideal-MHD value by a factor of $\sim 1.5$ over $R/R_{0}\sim 1$$10$ when the background anisotropy is perpendicular-temperature dominated. Moreover, in case 3, temperature anisotropy with $T_{\parallel}>T_{\perp}$ reduces the maximum growth rate, thereby limiting the radial extent of the PDI-unstable region to $R\lesssim 30R_{0}$. Other effects—such as solar-wind expansion, background turbulence, and Landau damping—are not considered in this study, so our estimated radial range for PDI occurrence could be confined to even smaller heliocentric distances. Our result demonstrates that anisotropy-driven enhancement remains effective in the parameter regime most relevant to PSP observations, and therefore temperature anisotropy should be treated as a non-negligible control parameter when assessing where PDI can operate in the near-Sun wind.
@@
    In the adiabatic case, the radial evolution of $\gamma_{\max}/\omega_{0}$ differs qualitatively between the isotropic and anisotropic closures. Under the adiabatic assumption, ideal MHD predicts that $\gamma_{\max}/\omega_{0}$ increases with heliocentric distance in the low-$\beta$ regime, with an approximate power-law scaling that approaches the classical ${\gamma}_{\max}/\omega_{0}\propto R^{1/3}$ behavior in the smaller-amplitude limit. In contrast, the CGL model does not exhibit a monotonic increase with $R$. Instead, ${\gamma}_{\max}/\omega_{0}$ can turn over and decrease beyond a certain distance. The parameter scans in Section IV.1 suggest that in the low-$\beta$ regime, the PDI growth rates in both models are nearly identical even when $\xi=10$ is imposed, whereas the direct $\xi$-dependence itself remains weak over a broad range. Because the adiabatic closures imply different radial scalings, $\beta\propto R^{2/3}$ in MHD and $\beta_{\parallel}\propto R^{2}$ in CGL, the more rapid increase of $\beta_{\parallel}$ in the CGL closure provides a plausible explanation for the different radial behaviors found in case 1.
@@
    Perturbed temperature anisotropy reduces ${\gamma}_{\max}/\omega_{0}$. Even when the background anisotropy is set to unity ($\xi=1$), the CGL-based dispersion relation does not reduce to the ideal-MHD form because the linearized pressure responses differ unless one further imposes $\delta T_{\perp}=\delta T_{\parallel}$ (for $n=5/3$). By comparing ${\gamma}_{\max}/\omega_{0}$ from CGL at $\xi=1$ with the ideal-MHD result, we quantified the impact of the perturbed temperature anisotropy. The resulting maps show that ${\gamma}_{\max}/\omega_{0}$ is generally smaller in CGL than in MHD, and the difference grows with increasing $\hat{B}_{\perp}$ and $\beta$. This implies that allowing $\delta T_{\perp}\neq\delta T_{\parallel}$ tends to suppress PDI.
These results have direct implications for PDI-based interpretations and models of the near-Sun solar wind. Because PDI produces compressive fluctuations and counter-propagating Alfvénic components, any temperature-anisotropy-driven modulation of ${\gamma}_{\max}/\omega_{0}$ can influence the generation of density fluctuations, the effective reflection rate, and the subsequent development of Alfvén-wave turbulence. Our findings therefore motivate incorporating anisotropic thermodynamics when estimating PDI activity in inner-heliosphere applications, including global simulations that use PDI as an effective mechanism for controlling turbulence.